# Bayesian-Driven Graph Reasoning for Active Radio Map Construction

### v1

@@ II-A Sampling Model
The agent traverses an ordered sequence of waypoints $\rho\triangleq\bigl{\langle}v_{1},v_{2},\dots,v_{n}\bigr{\rangle}$, with fixed starting point and ending point, i.e., $v_{1}=s$ and $v_{n}=g$. Let =$\mathcal{F}=\{(x,y)\mid\mathcal{M}(x,y)=0\}$ denotes the set of obstacle-free positions, then the $i$-th point within $\rho$, denoted as $\rho_{i}$ must belong to $\mathcal{F}$. The agent moves along straight-line segments connecting consecutive waypoints, ensuring that each segment remaining entirely within $\mathcal{F}$.
@@ III-B Uncertainty-Aware Graph-Based Path Planning
    The first case rewards the agent for reducing uncertainty from step $m-1$ to $m$, as long as the current node location $v_{m}$ is not the goal node location $g$. The second case applies a penalty proportional to the uncertainty $\Sigma^{(m)}$ if the agent reaches the goal node location $g$, with $\beta$ being a penalty factor. This setup aims to strategically guide the agent’s exploration by balancing the reduction of uncertainty with resource constraints, ultimately enhancing the efficiency and effectiveness of the exploration process.
@@ III-C Proximal Policy Optimization with Action Masking
The construction of this mask accounts for the residual budget $B_{m}$, collision checks, and a return-to-goal constraint, prunes infeasible neighbors before the policy is sampled. Neighbors deemed infeasible receive mask value 0; feasible ones receive 111. This effectively prunes infeasible neighbors before the policy is sampled.
@@
Let $\theta$ and $\theta_{\text{old}}$ denote the current and previous policy parameters, respectively. Define the likelihood ratio as $r_{m}(\theta)={\pi_{\theta}^{\text{masked}}(a_{m}\mid s_{m})}/{\pi_{\theta_{\text{old}}}^{\text{masked}}(a_{m}\mid s_{m})}$ To stabilize training, we compute the advantage denoted $\hat{A}_{m}$ using Generalized Advantage Estimation (GAE). GAE leverages a learned state-value function $V(s)$, provided by a critic network trained jointly with the policy.
@@
Epochs 500500500
Dropout prob. 0.10.10.1
RL Batch size 204820482048
@@ III-D Prediction and Planning Cycle
[…]
@@ IV Experiment
[…]
@@ IV-A Uncertainty-Aware Network
To train the radio map prediction network, we downsample all radio maps in the RadioMapSeer dataset to a spatial resolution of $\times 64$ 64. The entire dataset is split into training, validation, and test sets in a 7:2:1 ratio.

### v2

@@ II-A Sampling Model
The agent traverses an ordered sequence of waypoints $\rho\triangleq\bigl{\langle}v_{1},v_{2},\dots,v_{n}\bigr{\rangle}$, with fixed starting point and ending point, i.e., $v_{1}=s$ and $v_{n}=g$. Let $\mathcal{F}=\{(x,y)\mid\mathcal{M}(x,y)=0\}$ denotes the set of obstacle-free positions, then the $i$-th point within $\rho$, denoted as $\rho_{i}$ must belong to $\mathcal{F}$. The agent moves along straight-line segments connecting consecutive waypoints, ensuring that each segment remaining entirely within $\mathcal{F}$.
@@ III-B Uncertainty-Aware Graph-Based Path Planning
    The first case rewards the agent for reducing uncertainty from step m1m-1 to $m$, as long as the current node location $v_{m}$ is not the goal node location $g$. The second case applies a penalty proportional to the uncertainty $\Sigma^{(m)}$ if the agent reaches the goal node location $g$, with $\beta$ being a penalty factor. This setup aims to strategically guide the agent’s exploration by balancing the reduction of uncertainty with resource constraints, ultimately enhancing the efficiency and effectiveness of the exploration process.
@@ III-C Proximal Policy Optimization with Action Masking
The construction of this mask accounts for the residual budget $B_{m}$, collision checks, and a return-to-goal constraint, prunes infeasible neighbors before the policy is sampled. Neighbors deemed infeasible receive mask value 0; feasible ones receive $1$. This effectively prunes infeasible neighbors before the policy is sampled.
@@
Let $\theta$ and $\theta_{\text{old}}$ denote the current and previous policy parameters, respectively. Define the likelihood ratio as $r_{m}(\theta)={\pi_{\theta}^{\text{masked}}(a_{m}\mid s_{m})}/{\pi_{\theta_{\text{old}}}^{\text{masked}}(a_{m}\mid s_{m})}$ . To stabilize training, we compute the advantage denoted $\hat{A}_{m}$ using Generalized Advantage Estimation (GAE). GAE leverages a learned state-value function $V(s)$, provided by a critic network trained jointly with the policy.
@@
Epochs $500$
Dropout prob. $0.1$
RL Batch size $2048$
@@ III-D Prediction and Planning Cycle
The runtime of each URAM decision step is dominated by two main components: (i) the uncertainty-aware map prediction, which requires $K$ forward passes of the U-Net on an $N\times N$ map, leading to a complexity of $O(KN^{2})$; and (ii) the path planner, whose self-attention mechanism over $|\mathcal{V}|$ graph nodes results in a complexity of $O(|\mathcal{V}|^{2})$. The total complexity is therefore $O(KN^{2}+|\mathcal{V}|^{2})$. Through empirical evaluation, we determined that setting $K=10$ provides a robust and stable uncertainty map suitable for effective path planning, while avoiding the diminishing returns and higher computational burden of larger values. In contrast, reconstruction network like DRUE [spectrumsurveying] operates at a faster $O(N^{2})$ but it cannot quantify epistemic uncertainty. On the other hand, traditional methods like GPs provide exact variance but at an intractable cost of $O(N_{s}^{2}+N_{s}N^{2})$, where $N_{s}$ is the number of samples. This makes GPs impractical for real-time applications with large datasets ($N>10^{3}$). URAM with a small constant $(K=10)$, thus provides a superior balance, delivering robust uncertainty estimates with computational costs that remain competitive with deterministic methods and significantly more scalable than GPs.
@@ IV Experiment
To validate the URAM framework, all experiments are conducted on the IRT2 subset of the RadioMapSeer dataset [radiomapseer], a second-order Intelligent Ray Tracing (IRT) corpus that emulates realistic multipath propagation in urban settings. IRT2 comprises 700 distinct 256 m×256 $\text{ m}\times 256\text${ m} scenarios with heterogeneous building layouts and densities; each scenario is paired with 80 transmitter configurations, and the RSS is tabulated on a dense grid covering the entire area. From these 56000 scene–transmitter pairs we carve out 100 held-out environments—ten representative scenarios, each with ten transmitters—for policy training and benchmarking. This split guarantees diverse yet controlled test conditions for both radio-map reconstruction and uncertainty-aware path planning. The Root Mean Square Error (RMSE) is adopted as the error metric.
@@ IV-A Uncertainty-Aware Network
To train the radio map prediction network, we downsample all radio maps in the RadioMapSeer dataset to a spatial resolution of $64\times 64$. The entire dataset is split into training, validation, and test sets in a 7:2:1 ratio.